Trail Snax Corp. mixes raisins that cost $5.00 per kg with peanuts that cost $3.80 a kg. How many kilograms of raisins should be mixed with 10kg of peanuts to obtain a mixture worth $4.00 per kg?
Answer by stanbon(75887)

Trail Snax Corp. mixes raisins that cost $5.00 per kg with peanuts that cost $3.80 a kg. How many kilograms of raisins should be mixed with 10kg of peanuts to obtain a mixture worth $4.00 per
----
Equations:
Value Equation: 5r + 3.8*10 = 4(r + 10)
5r + 38 = 4r + 40
----
r = 2 kg (amt. of raisins needed in the mixture)
=====================
Cheers,
Stan H.